You have a patient that weighs 150 pounds. He has just been resuscitated and now has a BP of 60/40. You are going to establish a dopamine drip IV piggy back. You are using 60gtt set and wish to deliver 5 mcg/kg/min. You mix 200mg of dopamine into 250ml of normal saline. What is your flow rate?
Answer #1

ANSWER.

* Calculation of flow rate.

       Flow rate(gtt/min) = Volume(ml) / Time(min) * Drop factor(gtt/ml).

    Calculation of volume(ml).

         Volume(ml) = Desired dose/Dose in hand * Quantity.

      Here

          Ordered dose = 5 mcg/kg/min.

      Convert mcg to mg for calculation

              1 mcg = 0.001 mg.

       So

         Ordered dose = 5 * 0.001

                                 = 0.005 mg.

          Weight of the patient = 150 pound.

     Convert lbs to kg for calculation

                  1 lbs = 0.454 kg

         Weight of the patient = 150 * 0.454

                                           = 68 kg

            Desired dose(D) = 0.005 * 68

                                        = 0.34 mg.

           Dose in hand (H) = 200 mg.

                 Quantity (Q) = 250 ml

     As per above formula

             Volume(ml) = D/H * Q

                                = 0.34 / 200 * 250

                                = 0.0017 * 250

                                = 0.4 ml.

           Time(min) = 1 min.

             Drop factor = 60 gtt/ml.

     Flow rate = volume/Time * drop factor.

                     = 0.4 / 1 * 60

                     = 0.4 * 60

                     = 24 gtt/min.